Two cars leave towns 600 kilometers apart at the same time and travel toward each other. One car's rate is 18 kilometers per hour more than the other's. If they meet in 3 hours, what is the rate of the faster car?
Answer by mananth(16946)

The time taken by both the cars is same = 3 hours
let speed of one be x kmph
other's speed is (x+18) kmph
3x+3(x+18) = 600
3x+3x+54=600
6x+54=600
6x= 546
x= 546/6
x= 91 km/h
x+18 = 91+18 = 109 km/h ------------- second car